t**3 - 2 t**2 -48 t = 0
Answer by algebrahouse.com(1659)

t³ - 2t² - 48t = 0
t(t² - 2t - 48) = 0 {factored t out}
t(t - 8)(t + 6) = 0 {factored into two binomials}
t = 0 or t - 8 = 0 or t + 6 = 0 {set each factor equal to 0}
t = 0 or t = 8 or t = -6 {solved each equation for t}
www.algebrahouse.com